# Importance-based Token Merging for Diffusion Models

### v1

@@ Abstract
Diffusion models excel at high-quality image and video generation. However, a major drawback is their high latency. A simple yet powerful way to speed them up is by merging similar tokens for faster computation, though this can result in some quality loss. In this paper, we demonstrate that preserving important tokens during merging significantly improves sample quality. Notably, the importance of each token can be reliably determined using the classifier-free guidance magnitude, as this measure is strongly correlated with the conditioning input and corresponds to output fidelity. Since classifier-free guidance incurs no additional computational cost or requires extra modules, our method can be easily integrated into most diffusion-based frameworks. Experiments show that our approach significantly outperforms the baseline across various applications, including text-to-image synthesis, multi-view image generation, and video generation.
@@ 1 Introduction
Diffusion models [26, 89] excel in image and video generation [78, 79, 2, 17]. The rise of powerful diffusion models such as DALL-E [74], Stable Diffusion (SD) [78], or Imagen [79] has dramatically changed the landscape of generative AI. At their core, these models operate by iteratively denoising images or videos through multiple passes of a backbone network, processing a substantial number of tokens, which makes them particularly computationally intensive. For example, it could take up to 6 seconds to generate a single image with the Stable Diffusion model. A simple but powerful approach to speed up is token merging [4, 3] (ToMe), which merges similar or redundant tokens before executing computationally heavy operations, such as attention, making it highly effective for generating high-resolution images or videos.
ToMe frameworks typically divide tokens into two sets: source (src) and destination (dst) where most similar tokens from the src set are merged into the dst set. Essentially, tokens in the dst set function as anchors for grouping related tokens and also serve as the primary regions where computation will take place. Thus, token partitioning is particularly crucial to minimize performance drop and avoid image artifacts. To this end, ToMeSD [3] proposes to select a destination token randomly within each $2\times 2$ region, shown to be better than both fully random sampling and other fixed-pattern methods. However, this selection scheme is still random in nature and does not prioritize tokens based on relevance, which potentially misallocates resources to non-essential areas, as can be seen in Fig. 1 (a).
Wise old man reading a book by candlelightRefer to caption (a) ToMeSD [3] (b) Ours
@@
To address these limitations, we propose a novel token selection method that focuses computational resources on areas of high relevance to the prompt. Unlike the random selection used in ToMeSD [3], our approach leverages classifier-free guidance [25] as a targeted mechanism for token selection. While classifier-free guidance has been typically used as a global adjustment to emphasize prompt-relevant features, here we show that we can use it at the token level to dynamically highlight tokens carrying more essential information. By prioritizing token selection based on this prompt relevance, we ensure that high-information regions receive focused attention, preserving critical details and enhancing image quality. Further, classifier-free guidance incurs no additional computational cost and requires no modification to the architecture, thus, it can be readily used for any conditional diffusion models.
More specifically, our method assigns each token a relevance score based on its magnitude in the classifier-free guidance map, indicating its importance to the prompt. We then construct a pool of important tokens and only select destination tokens from this pool. Unlike simply picking the top-k important tokens, which can lead to redundancy and a suboptimal allocation of computing resources, our approach ensures a balance between relevance and diversity, focusing on prompt-aligned areas while maintaining efficiency. We apply our token merging strategy across three key applications: text-to-image generation, text-to-video generation, and multi-view image generation. In all cases, our method can be integrated seamlessly without requiring any modifications. Compared to the baseline method, our approach consistently demonstrates superior performance, delivering results with higher fidelity and significantly improved image details across all tested scenarios.
@@
  • We propose using classifier-free guidance magnitude as a measure of token importance. This can be obtained efficiently and robustly without any modification to diffusion models.
  • We design a novel token-merging strategy based on token importance to improve the quality of generated samples. We introduce an adaptive important token pool to enhance resource allocation.
  • Our method achieves state-of-the-art performance across various diffusion model tasks, including text-to-image synthesis, multi-view generation, and video generation.
@@ 2.1 Diffusion Models
Diffusion models [87, 12, 26, 89] are a class of generative models that iteratively transform random noise into complex data structures, such as images, by gradually reversing a diffusion process. In these models, data is progressively corrupted by adding noise, and the model learns to recover the original data through iterative denoising. This approach has demonstrated impressive results in generating high-quality, realistic images [78, 79, 70, 6, 74, 66, 16, 64, 104, 1], videos [83, 2, 17, 113, 38, 5], 3D content [72, 51, 56, 82, 41, 59, 67], and audio [36, 27].
@@
To reduce the cost of diffusion inferences, various techniques have been proposed. Some approaches require retraining, including better model structures [78, 71, 34], model pruning [14], model compression [112, 105, 46], step distillation [80, 65, 81, 52, 18], and consistency regularization [90, 60]. Others avoid retraining, such as improved sampling to reduce inference steps [89, 50, 57, 58], caching [62, 100, 8, 111, 31, 85, 61], model quantization [44, 7, 22, 94, 11, 86], and token reduction [32, 3, 4]. In this work, we focus on improving token reduction. Notably, our method is compatible with other diffusion acceleration techniques, enabling combined use to further improve inference efficiency.
@@ 2.2 Token Reduction
Recent studies have applied token reduction to diffusion model inference [3, 35, 21, 93, 45, 32, 110, 47, 95, 84]. ToMeSD [3] introduced token merging to Stable Diffusion [78] with a training-free method. It selects destination tokens and uses bipartite soft matching to merge redundant tokens. The reduced token set is processed by operations like attention, and then copied back to merged token locations to ensure the final token count matches the input. While ToMeSD performs well, its selection of destination tokens is not optimal, i.e., random across the image. In our work, we propose that selecting destination tokens based on their importance improves the fidelity and quality in generation, and this importance can be easily obtained via classifier-free guidance without additional cost. AT-EDM [93] uses self-attention maps to derive token importance, but this requires first performing full self-attention and increases peak memory usage. ATC [21] applies bottom-up hierarchical clustering for better token merging, but our experiment shows that this clustering is costly for generative tasks.
3 Preliminaries
Diffusion models.
Diffusion models [12, 26, 89] are generative models that produce data by simulating a denoising process. Starting with data $\mathbf{x}_{0}$, the forward diffusion process gradually adds Gaussian noise over $T$ timesteps, producing a sequence $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$. This process is defined as:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\,\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $\beta_{t}$ controls the noise schedule. The reverse process denoises the input by modeling $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ with learned parameters. Most models predict the noise added at each timestep, $\epsilon_{\theta}(\mathbf{x}_{t},t)$.
Classifier-free guidance (CFG) [25] modifies the noise prediction to improve sample control. It is designed to steer the predicted noise in a direction more aligned with the condition:
@@ 3 Preliminaries
where $\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)$ is the noise prediction conditioned on input $y$, $\epsilon_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction, and $w$ is the guidance weight.
Token Merging for Diffusion Models.
Our token merging method builds on ToMeSD [3], which is designed to speed up text-to-image diffusion models. To accelerate a model layer, such as attention, ToMeSD divides the input feature tokens into a source (src) set and a destination (dst) set. The dst set is selected randomly, i.e., by choosing one token from each $2\times 2$ region in the input feature map. Using bipartite soft matching, it then selects and merges the $r$ most similar tokens from the src set into the dst set, where $r$ represents the token merging ratio. The merging is done via averaging. After the model layer processes this reduced token set, the merged locations are filled with the corresponding processed dst tokens, maintaining the same output shape as the input.
4 Method
Figure 2: Overview. We propose an importance-based token merging method. The importance of each token is determined by the magnitude of classifier-free guidance. These scores, visualized with colors ranging from light to dark (indicating less to more important tokens), are used to construct a pool of important tokens. We randomly select a set of destination (dst) tokens from this pool and the remaining important tokens become source (src) tokens. Bipartite soft matching is then performed between the dst tokens and src tokens. src tokens without a suitable match are considered independent tokens (ind.). All other src tokens and unimportant tokens are merged with the destination tokens for subsequent computational steps.
We propose a novel token-merging method for diffusion models. Our approach uses classifier-free guidance to identify important tokens, allowing for a more efficient allocation of computing resources. Additionally, we create a dynamic pool of important tokens, with the pool size adapting to the token merging ratio. This adjustment enhances resource allocation and reduces redundancy when determining which tokens to merge. Fig. 2 provides an overview of our method.
Important Tokens.
Selecting which tokens to serve as anchors (destination tokens) for merging is critical for generating high-quality content. This is because these tokens correspond to the primary image regions where subsequent computations, such as attention, are applied. However, ToMeSD [3] currently relies on a random selection process based on a simple heuristic. It lacks well-established criteria for determining token importance. The primary guiding principle of ToMeSD is to prioritize spatial uniformity among the destination tokens [3, 4].
@@ 4 Method
Figure 3: Importance Maps. We show token importance maps derived from classifier-free guidance across diffusion inference timesteps. The generated image from the diffusion model is displayed on the left for reference.
In this paper, we highlight that classifier-free guidance [25] serves as a reliable indicator of token importance, providing a principled approach to identifying image regions with essential information. By leveraging this connection, we propose a method that prioritizes computing resources on important tokens, enabling more efficient and informed token merging even under constrained compute budgets. For each token, we calculate its importance as the absolute value of its classifier-free guidance score:
$\text{importance}=\left|\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)\right|,$ (3)
where $\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)$ is the noise prediction conditioned on input $y$ and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction. In Fig. 3, we provide an example of the importance maps.
A car made out of vegetablesRefer to caption (a) Top-k (b) Ours
Figure 4: We compare our method with an approach that uses the top-k important tokens as destination tokens (dst) for token merging. The computation locations after token merging are illustrated as non-black pixels in the bottom-left windows. They include locations of dst tokens, which are shown in white, and independent tokens (some other tokens that lack a similar dst token for merging), which are shown in orange.
Figure 5: Qualitative comparison of text-to-image generation. The first row shows results from Stable Diffusion (SD) [78], while the subsequent rows show SD combined with various token merging methods including ATC [21], ToMeSD [3], and our proposed method. Our approach consistently produces finer details with coherent structures. Note that ATC requires minutes to generate an image, whereas other methods, including ours, complete the task in seconds. The token merging ratio is 0.7. Please refer to the supplementary for detailed prompts. Best viewed with zoom-in for clarity.
Figure 6: Qualitative comparison of multi-view diffusion. We apply ToMeSD [3] and our token merging method to the multi-view diffusion model. We use Zero123++ [82] as the base model and a merging ratio of 0.6. Best viewed with zoom-in. Please refer to supplementary for input images.
@@
With the token importance scores, a naive approach is to pick the top-k tokens as destination tokens and merge the rest tokens that are similar to them. However, this approach produces low-quality outputs due to merging inefficiency, as shown in Fig. 4 (a). More specifically, there are two particular issues about this:
@@
    Unimportant Independent Tokens. In the standard token merging pipeline [3], there are “independent” tokens that don’t have similar destination tokens, which are processed independently without merging. In our case, background or irrelevant tokens can easily become independent tokens as there are no good matches for them among the important tokens, as shown in Fig. 4.
To this end, we propose to first create a pool of the most important tokens and ensure both destination tokens and independent tokens are drawn from this set. Destination tokens are randomly sampled from the pool, while independent tokens are selected as those in the important token pool that are most dissimilar to the chosen destination tokens. This simple method ensures that all subsequent computations after the merging step are performed on important tokens, leading to improved detail preservation, as demonstrated in Fig. 4 (b).
@@
More specifically, at time-step $t$ of the diffusion inference, a diffusion model layer, e.g. a transformer layer, takes $\mathbf{N}$ tokens as input. The token merging ratio is $r$. Based on the token importance derived from the previous timestep’s classifier-free guidance, we select the top $\mathbf{K}=\mathbf{N}\cdot(1-r)\cdot(1+p)$ tokens as the important token pool, denoted as $\mathbf{A}$. From $\mathbf{A}$, we randomly pick $\mathbf{D}=\mathbf{N}\cdot k$ tokens to form the destination (dst) set. Here, $p,k$ are the two hyper-parameters of our method. The remaining tokens in $\mathbf{A}$ become the source (src) set.
Next, we perform bipartite soft matching by computing pairwise cosine similarities between src and dst tokens. Each src token is linked to its most similar dst. Then, in src set, we select the top $\mathbf{I}=\mathbf{N}\cdot(1-k-r)$ tokens with the smallest similarity to their closest dst tokens to serve as independent tokens. The remaining src tokens and unimportant tokens are merged into their corresponding dst tokens. After merging, the number of tokens reduces to $\mathbf{I}+\mathbf{D}=\mathbf{N}\cdot(1-r)$ tokens.
Preserving Structure in Early Time-steps.
Prior work [35] suggests that token pruning (directly dropping tokens) could help preserve structural details. Building on this observation, we found that incorporate token pruning during the early diffusion steps, followed by token merging in later steps, improves generation results. Furthermore, the low token variance in the early steps [93] reduces the effectiveness of classifier-free guidance in identifying important tokens. To mitigate this, we simply employ the ToMeSD strategy during these initial steps, e.g., the first 6 out of 50 steps, to select one token from each $2\times 2$ region of the feature map as a destination token.

### v2

# Importance-Based Token Merging for Efficient Image and Video Generation

@@ Abstract
Token merging can effectively accelerate various vision systems by processing groups of similar tokens only once and sharing the results across them. However, existing token grouping methods are often ad hoc and random, disregarding the actual content of the samples. We show that preserving high-information tokens during merging—those essential for semantic fidelity and structural details—significantly improves sample quality, producing finer details and more coherent, realistic generations. Despite being simple and intuitive, this approach remains underexplored.
To do so, we propose an importance-based token merging method that prioritizes the most critical tokens in computational resource allocation, leveraging readily available importance scores, such as those from classifier-free guidance in diffusion models. Experiments show that our approach significantly outperforms baseline methods across multiple applications, including text-to-image synthesis, multi-view image generation, and video generation with various model architectures such as Stable Diffusion, Zero123++, AnimateDiff, or PixArt-$\alpha$.
@@ 1 Introduction
The rise of powerful diffusion models such as DALL-E [79], Stable Diffusion (SD) [83], or Imagen [84] has dramatically changed the landscape of generative AI [27, 94, 83, 84, 2, 18]. At their core, these models operate by iteratively denoising through multiple passes of a backbone network, processing a substantial number of tokens, which are particularly computationally intensive. To reduce the computational demands, prior work [4, 3, 37] has explored merging similar tokens and sharing the results across member tokens.
Prompt: “A cute owl wearing a wizard hat.Refer to caption (a) Spatial [3](b) Importance-based (Ours)
@@
However, such a merging process can also significantly degrade image quality, losing important details and structure. This typically happens when merging occurs in highly informative image regions, where critical visual details are compressed or discarded. One major reason is that existing methods rely on ad hoc heuristics for grouping tokens, often based on spatial proximity or fixed patterns without considering their content. Thus, destination tokens, where the other tokens will merge into, are chosen randomly within predefined regions, sometimes pushing important visual elements into less relevant areas. This leads to suboptimal resource allocation, where crucial textures, edges, or fine-grained structures are lost, while less important regions retain more detail than necessary, as can be seen in Fig. 1 (a). Existing methods rarely consider token content explicitly when merging.
To address this limitation, we propose a novel token selection method that prioritizes computational resources in areas of high visual and semantic importance. Unlike existing approaches that rely on random or fixed-pattern selection, our method ensures that merging decisions are guided by actual content relevance. Rather than treating all tokens equally, we use per-token importance signals to identify key regions and ensure that merging prioritizes preserving more important tokens. Note that there are many proxies for importance, such as attention scores, saliency maps, or user-provided bounding boxes, all of which can be integrated into our merging method. Specifically in our case, we point out that an excellent choice is classifier-free guidance (CFG) [26], as it inherently highlights regions that strongly influence model outputs and is obtainable with no additional cost.
More specifically, instead of merging tokens arbitrarily across the image, we first construct a pool of high-importance tokens. Then, token partitioning is only performed within this pool using a soft-matching strategy. This ensures that the most relevant tokens serve as anchors for merging, preserving key details while also maintaining diversity among the anchor tokens. Doing so facilitates a more efficient allocation of computational resources, preventing redundancy and ensuring that merging decisions are both meaningful and effective. Compared to simply selecting the top-k important tokens—which can lead to redundancy and suboptimal token assignments—our method strikes a balance between relevance and diversity, resulting in more coherent and detailed generations.
We apply our token merging strategy across three key applications: text-to-image generation, multi-view generation, and text-to-video generation, as well as across various model structures such as U-Net or transformers. Compared to baseline methods, our method consistently demonstrates superior performance, delivering results with higher fidelity and significantly improved image details across all tested scenarios.
@@
  • We propose a novel importance-based token merging paradigm for diffusion models, designed to preserve crucial image content. The importance score can be easily obtained from CFG at no additional cost.
  • We design a novel token-partitioning strategy based on a pool of important tokens to improve generation quality.
  • Our method achieves state-of-the-art performance across various diffusion model tasks, including text-to-image synthesis, multi-view generation, and video generation, as well as across model architectures including U-Net and transformers.
@@ 2.1 Diffusion Models
Diffusion models [92, 13, 27, 94] are a class of generative models that iteratively transform random noise into complex data structures, such as images, by gradually reversing a diffusion process. In these models, data is progressively corrupted by adding noise, and the model learns to recover the original data through iterative denoising with learned parameters. This approach has demonstrated impressive results in generating high-quality, realistic images [83, 84, 75, 7, 79, 71, 17, 69, 109, 1, 34], videos [88, 2, 18, 118, 40, 6], 3D content [77, 54, 59, 87, 43, 63, 72], and audio [38, 28]. A popular diffusion model architecture is introduced by Stable Diffusion [83], which uses a U-Net with transformer blocks. It first encodes a noisy image as latent tokens, which are processed through transformer blocks comprising self-attention, MLP, and cross-attention layers. This design has been extended for multi-view generation with multi-view attention [87] and for video generation with temporal layers [18]. More recent approaches [7, 118] replace the U-Net with a fully transformer-based architecture for improved scalability.
@@
To reduce the cost of diffusion inferences, various techniques have been proposed. Some approaches require retraining, including better model structures [83, 76, 36], model pruning [15], model compression [117, 110, 49], step distillation [85, 70, 86, 55, 19], and consistency regularization [95, 64]. Others avoid retraining, such as improved sampling to reduce inference steps [94, 53, 60, 61], caching [67, 105, 9, 116, 32, 90, 66, 45], model quantization [47, 8, 23, 99, 12, 91], and token reduction [33, 3, 4]. In this work, we focus on improving token reduction. Notably, our method is compatible with other diffusion acceleration techniques, enabling combined use for further speedup.
@@ 2.2 Token Reduction
Recent studies have applied token reduction to diffusion model inference [3, 37, 22, 98, 48, 33, 115, 50, 100, 89, 62]. ToMeSD [3] introduced token merging to Stable Diffusion [83] with a training-free method. The typical token merging procedure for diffusion models selects destination tokens from input feature tokens and utilizes bipartite soft matching to merge redundant tokens. The reduced token set is processed by operations like attention, and is then copied back to the merged token locations, ensuring the final token count matches the input. Although this token merging approach performs well, its selection of destination tokens is not optimal, i.e., spatially random across the image. ToFu [37] combines token pruning and merging but still follows similarly suboptimal destination token selection strategy. In our work, we propose that selecting destination tokens based on their importance improves the fidelity and quality in generation, and this importance can be easily obtained via classifier-free guidance without additional cost. AT-EDM [98] uses self-attention maps to derive token importance, but this requires first performing full self-attention and increases peak memory usage. ATC [22] applies bottom-up hierarchical clustering for better token merging, but we show it is costly for generative tasks.
3 Method
We propose a novel importance-based token-merging method, summarized in Fig. 2, allowing for a more efficient allocation of computing resources. We highlight that classifier-free guidance serves as an effective importance indicator. Further, we propose using a dynamic pool of important tokens, where the pool size adapts to the token merging ratio, optimizing resource allocation and reducing redundancy.
Figure 2: Overview. We propose an importance-based token merging method. The importance of each token can be determined using classifier-free guidance. These scores, visualized with colors ranging from light to dark (indicating less to more important tokens), are used to construct a pool of important tokens. We randomly select a set of destination (dst) tokens from this pool and the remaining important tokens become source (src) tokens. Bipartite soft matching is then performed between the dst tokens and src tokens. src tokens without a suitable match are considered independent tokens (ind.). All other src tokens and unimportant tokens are merged with the destination tokens for subsequent computational steps.
Important Tokens.
Selecting which tokens to serve as anchors (destination tokens) for merging is critical for generating high-quality content. This is because these tokens correspond to the primary image regions where subsequent computations, such as attention, are applied. In principle, all reliable per token importance signals, such as cross attention maps, user provided bounding boxes, or saliency maps, can be integrated with our method. We find that classifier-free guidance (CFG) [26] serves as an excellent indicator. CFG modifies the noise prediction to improve sample control. It is designed to steer the predicted noise in a direction more aligned with the condition:
@@ 3 Method
where $\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)$ is the noise prediction conditioned on input $y$, $\epsilon_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction, and $w$ is the guidance weight. It is widely used in diffusion models and incurs no additional computational cost. For each token, we calculate its importance as the absolute value of its CFG score:
importance $=\left|\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)\right|$ (2)
$\approx\left|-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p(y\mid\mathbf{x}_{t})\right|,$
where $\sigma_{t}$ is the noise scale. The guidance term effectively estimates the gradient of the log-likelihood of the conditioning variable $y$ (e.g., a text prompt) with respect to the noisy sample $\mathbf{x}_{t}$ [26]. Thus, the CFG magnitude can be interpreted as a saliency or importance measure: tokens with a high CFG magnitude have stronger influences in steering the generation toward satisfying the condition $y$. In Fig. 3, we provide an example of the resulting importance maps.
@@
Figure 3: Importance Maps. We present token importance maps derived from classifier-free guidance (CFG) across diffusion inference timesteps. These maps highlight areas significantly align with the user prompt. In the early steps, they capture the semantics and structure of the image relevant to the prompt, while in later steps, they focus on finer details of the objects the user intends to generate. The generated image is shown on the left for reference.
@@
With the token importance scores, a naive approach is to pick the top-k tokens as destination tokens (dst) and merge the rest tokens that are similar to them. However, this approach produces low-quality outputs due to merging inefficiency, as shown in Fig. 4 (a). More specifically, there are two particular issues about this:
@@
    Unimportant Independent Tokens. In the token merging pipeline, some tokens lack similar destination tokens, making them “independent” and remain unmerged. In the top-k approach, background or irrelevant tokens often become independent due to the absence of suitable matches among the important tokens, as shown in Fig. 4.
To avoid these issues, we propose to first create a pool of the most important tokens and then, ensure that both destination tokens and independent tokens are drawn from this set. To do so, destination tokens are randomly sampled from the pool, while independent tokens are selected as those in the important token pool that are most dissimilar to the chosen destination tokens. This approach ensures that all computations following the merging step operate on important tokens, leading to improved detail preservation, as illustrated in Fig. 4 (b).
@@
A delicate pink rose in full bloom, detailed petals.Refer to caption (a) Top-k(b) Ours
Figure 4: We compare our method with an approach that uses the top-k important tokens as destination tokens (dst) for token merging. The computation locations after token merging are illustrated as non-black pixels in the bottom-left windows. They include locations of dst tokens, which are shown in white, and independent tokens (some other tokens that lack a similar dst token for merging), which are shown in orange. Our method produces more structured and detailed image, as highlighted in the red box.
(a) SD [83] $\sim$ 8s(b) ATC [22] $\sim$ mins(c) ToFu [37] $\sim$ 5s(d) ToMe. [3] $\sim$ 5s(e) Ours $\sim$ 5s
Figure 5: Qualitative comparison of text-to-image generation. The first column shows results from Stable Diffusion (SD) [83], while the subsequent columns show SD combined with various token merging methods. As highlighted in red boxes, our approach consistently produces finer details with coherent structures. Note that ATC requires minutes to generate an image, whereas other methods, including ours, complete the task in seconds. The token merging ratio is 0.7. Please see the supplementary for prompts. Best viewed with zoom-in.
@@
At time-step $t$ of the diffusion inference, a diffusion model layer, e.g. a transformer layer, takes $\mathbf{N}$ tokens as input. The token merging ratio is $r$. Based on the token importance derived from the previous timestep’s classifier-free guidance, we select the top $\mathbf{K}=\mathbf{N}\cdot(1-r)\cdot(1+p)$ tokens as the important token pool, denoted as $\mathbf{A}$. From $\mathbf{A}$, we randomly pick $\mathbf{D}=\mathbf{N}\cdot k$ tokens to form the destination (dst) set. Here, $p$, $k$ are hyper-parameters. The remaining tokens in $\mathbf{A}$ become the source (src) set.
Next, we perform bipartite soft matching by computing pairwise cosine similarities between src and dst tokens. Each src token is linked to its most similar dst. Then, in src set, we select the top $\mathbf{I}=\mathbf{N}\cdot(1-k-r)$ tokens with the smallest similarity to their closest dst tokens to serve as independent tokens. The remaining src tokens and unimportant tokens are merged into their corresponding dst tokens. The merging is performed via averaging all grouped tokens. After merging, the number of tokens reduces to $\mathbf{I}+\mathbf{D}=\mathbf{N}\cdot(1-r)$ tokens. The diffusion model layer then processes this reduced token set, the merged locations are filled with the corresponding processed dst tokens, maintaining the same output shape as the input.